# DFML: Decentralized Federated Mutual Learning

### v1

@@ D.1 Convergence Speedup
D.2 Weighted vs Vanilla Average of KL Divergences (KLs)
In Equation (3), we use a weighted average of KL divergences (KLs) between all teacher models and the student. The weighting is determined based on the number of trainable parameters in each teacher model. Figure 15 demonstrates that the weighted average leads to better convergence speed and global accuracy compared to vanilla averaging. The reason is that larger models tend to have a higher probability of possessing finer knowledge, thus giving them more weight during knowledge distillation results in better knowledge transfer.
D.3 Effect of Increasing Mutual Learning Epochs $\boldsymbol{K}$
Figure 15 shows that increasing the number of mutual learning epochs $K$ at each aggregator contributes to a faster convergence speed. A higher number of $K$ enables more knowledge to be distilled from the experts, leading to improved convergence speed. This, in turn, results in a more efficient communication cost throughout training.
Figure 14: Effect of increasing mutual learning epochs $K$.
Figure 14: Effect of increasing mutual learning epochs $K$.
Figure 15: Evaluating DFML against decentralized with fewer number of senders $S$.
D.4 Different Number of Participants
@@
D.5 DFML vs Decentralized FML
In decentralized FML, the clients’ heterogeneous models are the same as the models used in our proposed DFML, which are based on CNN architectures and mode $H2$ from Table 7. The homogeneous model size used is $[32,64]$, which is the median of models of the five different architectures. Figure 16 compares our DFML and decentralized FML. We include the global performance of the local heterogeneous and the homogeneous meme models. We observe that the homogeneous meme models have a better convergence speed than our DFML, and in some cases, lead to better final accuracy. However, the global knowledge performance of heterogeneous models in decentralized FML is much worse than DFML and is even lower than decentralized FedAvg. As our goal is to transfer global knowledge to clients’ heterogeneous models, we consider that our DFML significantly outperforms the decentralized FML framework.
Appendix E Cyclical Knowledge Distillation: Further Analysis
E.1 Cyclical Knowledge Distillation
In Figure 17, we illustrate the impact of varying the value of $\alpha$ on global knowledge throughout training with full-cycle oscillations. Compared to Figure 5, it is evident that using a half-cycle oscillation results in better convergence speed. With half-cycle oscillations, less time is spent in the stage where the supervision signal is dominant, leading to less loss of global knowledge compared to full-cycle oscillations. When the distillation signal becomes dominant again, the recovery from the lost global knowledge is faster.
Figure 17: DFML performance against varying $\alpha$. The regular models are evaluated instead of the peak models and $\alpha$ oscillates with full cycles. The supervision signal used is CE.
Figure 18: Comparison between different supervision signals.
@@ E.1 Cyclical Knowledge Distillation
E.2 Different Supervision Signals
E.3 Different Adaptive Techniques
@@
E.4 Different Schedulers for Loss Components
In Equation (1), the objective function consists of two loss components: supervision and distillation loss signals. In Equation (9), we examine the impact of independently varying each loss component on global performance. Specifically, we compare additional scenarios where one scheduler is assigned to $\mathcal{L}_{\text{CE}}$ alone, another scenario where a scheduler is assigned to $\mathcal{L}_{\text{KL}}$ alone, and a third scenario where one scheduler scales both loss components with the same factor. For this experiment, we use CE as the supervision signal, as the improvements are more notable under the non-IID distribution shifts. Figure 20 illustrates the comparison of assigning independent schedulers to loss components. In this experiment, $\gamma$ oscillates from $1\rightarrow 0$, and $\alpha$ oscillates from $0\rightarrow 1$.
@@ E.4 Different Schedulers for Loss Components
Scaling the distillation signal alone and leaving the scale of $\mathcal{L}_{\text{CE}}$ untouched does not yield any advancement in performance compared to keeping $\alpha$ fixed for both loss components. This indicates that the supervision signal has a more significant impact than the distillation signal. Performance gains are observed when the $\mathcal{L}_{\text{CE}}$ signal is reduced, allowing more influence on the $\mathcal{L}_{\text{KL}}$ signal. In contrast, when $\mathcal{L}_{\text{CE}}$ oscillates while the $\mathcal{L}_{\text{KL}}$ scale is kept fixed, it results in the same performance as when both $\mathcal{L}_{\text{CE}}$ and $\mathcal{L}_{\text{KL}}$ signals are scaled in opposite directions (Equation (1)). This is because the $\mathcal{L}_{\text{CE}}$ signal is dominant without any scaling, and as it diminishes it allows the $\mathcal{L}_{\text{KL}}$ signal to take precedence. The peak value is attained when the $\mathcal{L}_{\text{CE}}$ signal reaches 00, and the $\mathcal{L}_{\text{KL}}$ signal’s scale is $1$. Finally, scaling both $\mathcal{L}_{\text{CE}}$ and $\mathcal{L}_{\text{KL}}$ signals with a common scheduler leads to inferior performance compared to scaling $\mathcal{L}_{\text{CE}}$ alone or scaling both signals in opposite directions. The poor performance of oscillating $\mathcal{L}_{\text{KL}}$ signal alone is attributed to the continuous dominance of $\mathcal{L}_{\text{CE}}$ signal during mutual learning, causing the models to drift toward the aggregator’s local data.
E.5 Fixed vs Increasing Period

### v2

@@ A.4.4 Convergence Speedup
A.4.5 Local Accuracy
Our proposed DFML not only surpasses the baselines in global accuracy but also achieves competitive results in local accuracy. As shown in Figure 18, the local accuracy attained by DFML generally exceeds that of decentralized FedAvg. Although DFML exhibits slightly lower local accuracy compared to decentralized FedAvg in the CIFAR-10 non-IID experiment with 10 clients, it remains competitive. Moreover, the corresponding global accuracy achieved by DFML in that experiment surpasses decentralized FedAvg.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 18: Local accuracy comparison between DFML and decentralized FedAvg for different numbers of clients and data distributions using non-restrictive heterogeneous architectures.
A.4.6 Different Number of Participants
@@
A.4.7 Weighted vs Vanilla Average of KL Divergences (KLs)
In Equation 3, we use a weighted average of KL divergences (KLs) between all teacher models and the student. The weighting is determined based on the number of trainable parameters in each teacher model. Figure 21 demonstrates that the weighted average leads to better convergence speed and global accuracy compared to vanilla averaging. The reason is that larger models tend to have a higher probability of possessing finer knowledge, thus giving them more weight during knowledge distillation results in better knowledge transfer.
A.4.8 Mutual vs Vanilla Knowledge Transfer
With vanilla knowledge transfer, only the aggregator’s model is updated, with the other models act as teachers. Conversely, with mutual learning, all models are updated. Figure 21 demonstrates the significant improvement in global accuracy and convergence speed achieved through mutual learning compared to vanilla knowledge transfer.
Figure 19: Evaluating DFML against decentralized FedAvg with fewer number of senders $S$.
Figure 21: Mutual vs vanilla knowledge transfer.
A.4.9 Effect of Increasing Mutual Learning Epochs $K$
Figure 22 shows that increasing the number of mutual learning epochs $K$ at each aggregator contributes to a faster convergence speed. A higher number of $K$ enables more knowledge to be distilled from the experts, leading to improved convergence speed. This, in turn, results in a more efficient communication cost throughout training. On the other hand, the computational cost will increase significantly. From the figure we can see that increasing the $K$ beyond 10 does not result in significant enhancement in convergence speed at the expense of adding more communications between clients. Further, selecting $K$ lower than 10 will save computation but will increase the communication cost. Thus, there is a trade-off between computation and communication cost for convergence. Last, Even though decentralize FedAvg is more computation efficient than DFML, but it does not reach to the same global accuracy level as DFML.
Figure 22: Effect of increasing mutual learning epochs $K$ on global accuracy with respect to (Left) communication rounds and (Right) computation cost.
A.4.10 Different Topology
In all previous experiments, we used a mesh topology, where all clients can reach each other. However, in this subsection, we explore a different network topology, as illustrated in Figure 23. In this new topology, clients are divided into two groups, each connected in a mesh configuration, with a single link connecting the two groups. The first group contains clients with odd addresses $[1,3,5,…,N-1]$ and the second group contains clients with even addresses $[2,4,6,…,N]$. The clients connecting group 1 and group 2 have the ”median” address from the list of addresses in its group in their respective groups.
We conducted experiments using both restrictive and nonrestrictive heterogeneous architectures. The results, provided in Tables 11 and 12, demonstrates that even with this different topology, DFML continues to surpass the baselines.
Figure 23: Different Topology.
Table 11: Global accuracy comparison using restrictive heterogeneous architectures with $50$ clients and $25$ senders. The experiments are conducted using ResNet architectures using the topology illustrated in Figure 23.
CIFAR-10 CIFAR-100
Method IID non-IID IID non-IID
Dec. FedAvg 44.71$\pm$ 0.12 22.11$\pm$ 0.06 10.36 $\pm$ 0.16 7.51 $\pm$ 0.14
Dec. HeteroFL 58.91 $\pm$ 0.17 43.45 $\pm$ 0.21 18.24 $\pm$ 0.13 14.81 $\pm$ 0.22
Dec. FedRolex 58.91 $\pm$ 0.09 43.54 $\pm$ 0.19 18.20 $\pm$ 0.08 15.10 $\pm$ 0.11
DFML (Ours) 78.87 $\pm$ 0.01 68.52 $\pm$ 0.02 45.02 $\pm$ 0.02 40.31 $\pm$ 0.03
Table 12: Global accuracy comparison using nonrestrictive heterogeneous architectures with $50$ clients and $25$ senders. The experiments are conducted using CNN architectures using the topology illustrated in Figure 23.
CIFAR-10 CIFAR-100
Method IID non-IID IID non-IID
Dec. FedAvg 51.45 $\pm$ 0.01 25.03 $\pm$ 0.02 27.10 $\pm$ 0.11 20.16$\pm$ 0.21
DFML (Ours) 78.87 $\pm$ 0.10 68.52 $\pm$ 0.18 45.01 $\pm$ 0.15 40.30$\pm$ 0.15
A.4.11 DFML vs Decentralized FML
In decentralized FML, the clients’ heterogeneous models are the same as the models used in our proposed DFML, which are based on CNN architectures and mode $H2$ from Table 10. The homogeneous model size used is $[32,64]$, which is the median of models of the five different architectures. Figure 24 compares our DFML and decentralized FML. We include the global accuracy of the local heterogeneous and the homogeneous meme models. We observe that the homogeneous meme models have a better convergence speed than our DFML, and in some cases, lead to better final accuracy. However, the global knowledge performance of heterogeneous models in decentralized FML is much worse than DFML and is even lower than decentralized FedAvg. As our goal is to transfer global knowledge to clients’ heterogeneous models, we consider that our DFML significantly outperforms the decentralized FML framework.
@@ A.4.11 DFML vs Decentralized FML
A.5 Cyclical Knowledge Distillation: Further Analysis
A.5.1 Different Supervision Signals
A.5.2 Different Adaptive Techniques
@@
A.5.3 Different Schedulers for Loss Components
In Equation 1, the objective function consists of two loss components: supervision and distillation loss signals. In Equation 10, we examine the impact of independently varying each loss component on global accuracy. Specifically, we compare additional scenarios where one scheduler is assigned to $\mathcal{L}_{\text{CE}}$ alone, another scenario where a scheduler is assigned to $\mathcal{L}_{\text{KL}}$ alone, and a third scenario where one scheduler scales both loss components with the same factor. For this experiment, we use CE as the supervision signal, as the improvements are more notable under the non-IID distribution shifts. Figure 27 illustrates the comparison of assigning independent schedulers to loss components. In this experiment, $\gamma$ oscillates from $1\rightarrow 0$, and $\alpha$ oscillates from $0\rightarrow 1$.
@@ A.5.3 Different Schedulers for Loss Components
Scaling the distillation signal alone and leaving the scale of $\mathcal{L}_{\text{CE}}$ untouched does not yield any advancement in performance compared to keeping $\alpha$ fixed for both loss components. This indicates that the supervision signal has a more significant impact than the distillation signal. Performance gains are observed when the $\mathcal{L}_{\text{CE}}$ signal is reduced, allowing more influence on the $\mathcal{L}_{\text{KL}}$ signal. In contrast, when $\mathcal{L}_{\text{CE}}$ oscillates while the $\mathcal{L}_{\text{KL}}$ scale is kept fixed, it results in the same performance as when both $\mathcal{L}_{\text{CE}}$ and $\mathcal{L}_{\text{KL}}$ signals are scaled in opposite directions (Equation 1). This is because the $\mathcal{L}_{\text{CE}}$ signal is dominant without any scaling, and as it diminishes it allows the $\mathcal{L}_{\text{KL}}$ signal to take precedence. The peak value is attained when the $\mathcal{L}_{\text{CE}}$ signal reaches 00, and the $\mathcal{L}_{\text{KL}}$ signal’s scale is $1$. Finally, scaling both $\mathcal{L}_{\text{CE}}$ and $\mathcal{L}_{\text{KL}}$ signals with a common scheduler leads to inferior performance compared to scaling $\mathcal{L}_{\text{CE}}$ alone or scaling both signals in opposite directions. The poor performance of oscillating $\mathcal{L}_{\text{KL}}$ signal alone is attributed to the continuous dominance of $\mathcal{L}_{\text{CE}}$ signal during mutual learning, causing the models to drift toward the aggregator’s local data.
A.5.4 Fixed vs Increasing Period